The manager of the Texas Department of Transportation has determined that it typically takes 30 minutes for the department's employees to register a new car. In Bexar County, the predetermined fixed overhead rate was computed on an estimated 10,000 direct labor hours per month and is $9 per direct labor hour, whereas the predetermined variable overhead rate is $3 per direct labor hour. During July, 18,800 cars were registered in Bexar County and 9,500 direct labor hours were worked in registering those vehicles. For the month, variable overhead was $27,700 and fixed overhead was $90,800.

A) Compute overhead variances using a four-variance approach.

VOH Spending Variance: $

VOH Efficiency Variance: $

Total VOH Variance: $

FOH Spending Variance $

FOH Volume Variance $

Total FOH Variance $

B) Compute overhead variances using a three-variance approach

OH Spending Variance $

OH Efficiency Variance $

OH Volume Variance $

Total OH Variance $

C) Compute overhead variances using a two-variance approach.

Budget Variance $

Volume Variance $

Total OH Variance $
Answer #1

Solution A:

Standard rate of VOH = $3

Actual rate of VOH = $27,700 / 9500 = 2.915789

Standard hours of VOH = 18800*0.5 = 9400 hours

Actual hours of VOH = 9500 hours

VOH Spending variance = (SR -AR) * AH = ($3 - $2,915789) * 9500 = $800 F

VOH Efficiency variance = (SH - AH) * SR = (9400 - 9500) * $3 = $300 U

Total VOH Variance = $800 F + $300 U = $500 F

Budgeted fixed overhead = 10000*9 = $90,000

Fixed overhead applied = 9400 * 9 = $84,600

Actual fixed overhead = $90,800

FOH Spending variance = Budgeted FOH - Actual FOH = $90,000 - $90,800 = $800 U

FOH Volume variance = FOH Applied - Budgeted FOH = $84,600 - $90,000 = $5,400 U

Total FOH Variance = $800 U + $5,400U = $6,200 U

Solution B:

OH spending variance = Budgeted overhead at actual hours - Actual overhead

= ($90,000 + 9500*$3) - ($27,700 + $90,800) = $0

OH efficiency variance = (SH - AH) * SR for VOH = (9400 - 9500)*$3 = $300 U

OH Volume variance = Overhead applied - Budgeted overhead

= (9400*$12) - ($90,000 + 9400*$3) = $5,400 U

Total OH Vairance = 0 + $300 U + $5,400 U = $5,700 U

Solution C:

Budget variance = Budgeted overhead - Actual overhead

= ($90,000 + 9400*$3) - ($90,800 + $27,700) = $300 U

Volume variance = Overhead applied - Budgeted overhead

= (9400*$12) - ($90,000 + 9400*$3) = $5,400 U

Total OH Vairance = $300 U + $5,400 U = $5,700 U